### v2

# DiffAdapt: Difficulty-Adaptive Reasoning for Token-Efficient LLM Inference

Abstract
Recent reasoning Large Language Models (LLMs) demonstrate remarkable problem-solving abilities but often generate long thinking traces whose utility is unclear. Our work aims to improve their efficiency, enabling them to reach high performance without overthinking. First, we analyze the entropy of token probabilities in reasoning traces. Across three models, we observe a consistent U-shaped entropy pattern: high entropy on easy problems despite high accuracy, low entropy on problems with medium difficulty, and high entropy on hard problems reflecting uncertainty. Specifically, we notice 22–25% entropy reduction from easy to medium difficulty regions, suggesting an overthinking phenomenon on easy instances. Building on these insights, we introduce DiffAdapt, a lightweight framework that selects Easy/Normal/Hard inference strategies per question based on their difficulty and reasoning trace entropy. Each inference strategy consists of a fixed prompt, temperature and maximum token length. In contrast to existing efficiency optimization methods, our approach does not fine-tune base LLM but a small probe that classifies LLM’s final hidden state, allowing inexpensive adaptation. We comprehensively evaluate our method on five models and eight benchmarks. Our method achieves comparable or improved accuracy while reducing token usage by up to 22.4%, establishing a practical path toward compute-efficient reasoning.
1 Introduction
Large Language Models (LLMs) emerged as powerful tools for complex reasoning tasks, spanning mathematical problem-solving (Lewkowycz et al., 2022a; Pan et al., 2024), code generation (Chen et al., 2021), and logical deduction (Wei et al., 2022). A key ingredient in this success is intermediate reasoning steps, often referred to as a “chain of thought” (CoT), before producing a final answer (Yu et al., 2024; Li et al., 2025d; Liu et al., 2024). However, this capability comes at a significant computational cost. Models typically generate a lengthy and elaborate chain of thought for every problem, a process sometimes called test-time scaling (Muennighoff et al., 2025; Chu et al., 2025).
Always generating long traces is fundamentally inefficient. It squanders resources on simple problems, while not necessarily providing sufficient resources for truly complex tasks (Qu et al., 2025; Sui et al., 2025; Li et al., 2025b). In this work, we introduce a framework to bridge this gap through systematic empirical analysis and adaptive inference strategy design. We first discover a U-shaped entropy pattern: high entropy on easy problems despite high accuracy, low entropy on problems with medium difficulty, and high entropy on hard problems reflecting uncertainty. Counterintuitively, models show high uncertainty on simple problems despite achieving high accuracy.
This motivates us to design a different inference strategy per each of this three distinct difficulty regions. We develop three simple inference strategies, each equipped with different generation length (i.e., max token length), prompt and decoding hyperparamenters (e.g., sampling temperature). The prompt designed for “easy” questions encourages models to answer succintly without overthinking, while the prompt for high difficulty questions instructs model to think carefully. We first conduct oracle experiments with these three templates. When allowed to choose an optimal strategy per question, models achieve 50% token savings while improving the model accuracy by over 10%.
Based on this observation, we introduce DiffAdapt, a three-stage framework that dynamically selects inference strategy rather than applying uniform reasoning budgets to all problems. Our framework operates in three stages: (1) we use a proxy model to generate training data by sampling responses and heuristically labeling them with difficulty-based strategy assignments; (2) we train a lightweight probe on the model’s hidden states to predict problem difficulty; and (3) during inference, the probe dynamically selects the appropriate reasoning strategy (Easy/Normal/Hard) for each question. Compared to the training-free baseline DEER (Yang et al., 2025b), DiffAdapt achieves superior performance with up to 62% token reduction and 18% performance improvement across eight mathematical reasoning benchmarks Lightman et al. (2023); Rein et al. (2024); Wang et al. (2024); He et al. (2024); Cobbe et al. (2021) over five models DeepSeek-AI et al. (2025); Yang et al. (2025a); Liu et al. (2025a); Hou et al. (2025).
2 Related Work
Training-Based Budget Control. Several methods incorporate budget control directly into the model’s training phase. Huang et al. (2025); Shen et al. (2025); Liu et al. (2025b) use reinforcement learning (RL) with difficulty-aware rewards to train a model for adaptive budgeting. Similarly, Cheng et al. (2025) employ a dual-reward system based on Group-Policy Optimization (GRPO) to encourage conciseness. ThinkPrune (Hou et al., 2025) trains long-thinking LLMs via RL with token limits, using iterative pruning rounds to achieve better performance-length tradeoffs. LC-R1 (Cheng et al., 2025) addresses “invalid thinking” through GRPO-based post-training with length and compress rewards, achieving significant sequence length reduction while maintaining performance. TL;DR (Li et al., 2025c) is a training-free method that uses mixed system1 and system2 data to control the reasoning process. AdaCoT (Lou et al., 2025) framed adaptive reasoning as a Pareto optimization problem that seeks to balance model performance with the costs associated with CoT invocation. Thinkless (Fang et al., 2025) is trained under a reinforcement learning paradigm and employs two control tokens, <short> for concise responses and <think> for detailed reasoning.
Inference-Time Budget Control. Other methods operate purely at inference time without requiring training. Zhang et al. (2025a) define an $\alpha$ moment” to switch from slow to fast thinking, while Zhang et al. (2025c) modify the sampling strategy to explore a continuous concept space. Yang et al. (2025b) monitor for specific transition tokens and model confidence to perform an early exit. Ma et al. (2025) propose a method to disable the reasoning process of LLMs. These training-free methods are flexible but often underperform a learned difficulty model.
Methods Requiring Auxiliary Models. Some approaches rely on external models to guide the LLM’s reasoning process. For instance, Li et al. (2025a) train a separate BERT model to predict the remaining reasoning length and steer the generation process. Zhang et al. (2025b) employ R1-7B as a switcher model, using prompt engineering or supervised fine-tuning for strategy selection. Liang et al. (2025) utilize an MLP-based switcher with group accuracy as training labels. While these methods can achieve good performance, they introduce the overhead of running additional non-trivial models during inference, increasing computational costs and deployment complexity. Our DiffAdapt framework, in contrast, is a very small classifier integrated directly with the LLM’s internal states, adding minimal latency.
3 Characterizing Overthinking Phenomenon
Recent work has identified that reasoning LLMs exhibit “overthinking” behavior (Ma et al., 2025; Sui et al., 2025; Qu et al., 2025), where models generate exceedingly lengthy solution when they can arrive at correct solutions much more succintly. Building upon this observation, we analyze this phenomenon from an entropy perspective, revealing a counterintuitive pattern where models demonstrate high uncertainty.
3.1 Experimental Setting
Dataset We use DeepMath-103K dataset (He et al., 2025), a large-scale, challenging, decontaminated, and verifiable mathematical dataset designed for advancing reasoning capabilities. The dataset provides problems with difficulty ratings from 1-10 as evaluated by GPT-4o based on mathematical complexity and covers a wide range of mathematical topics with rigorous decontamination against numerous benchmarks. For our experiments, we created a balanced experimental set by sampling 300 questions per difficulty level, with 10 sampling iterations per problem at temperature 0.6, ensuring robust statistical analysis of entropy patterns.
Figure 1: Visualization of model accuracy (blue bar), generation entropy (red line) per difficulty of question (x-axis), across three models (model name at the bottom of the graph). We observe a consistent U-shaped entropy curve along the difficulty levels on multiple models.
Entropy Calculation We measure model uncertainty using generation entropy, calculated as the average entropy across all tokens in the generated sequence (Wang et al., 2025). For each token position $t$, the entropy is computed as $H_{t}=-\sum_{j=1}^{V}p_{t,j}\log p_{t,j}$, where $V$ is the vocabulary size and $p_{t,j}$ is the probability of predicting the $j$-th token in the vocabulary given all preceding tokens $o_{<t}$. The entropy $H_{t}$ corresponds to the uncertainty of the token generation distribution at position $t$, rather than an intrinsic property of the specific token $o_{t}$ sampled from that distribution.
Correctness Rate Evaluation For each difficulty rating bucket, we measure the correctness rate as the fraction of successful solutions across multiple sampling iterations. For each problem $x_{i}$, we generate $n$ responses $\{r_{i,j}\}_{j=1}^{n}$ and evaluate their correctness against ground truth solutions. The correctness rate for problem $x_{i}$ is defined as:
$\mathcal{C}(x_{i})=\frac{1}{n}\sum_{j=1}^{n}\mathbb{I}[r_{i,j}=y_{i}]$
where $\mathbb{I}[\cdot]$ is the indicator function, $y_{i}$ denotes the ground truth solution for problem $x_{i}$, and $n=10$ sampling iterations per problem in our experiments.
3.2 Results
Figure 1 (a) illustrates this U-shaped entropy curve across different difficulty levels using the DeepSeek-R1-Distill-Qwen-7B model on the DeepMath-103K dataset (He et al., 2025). We observe consistent overthinking patterns across multiple model architectures (see Appendix D for additional models).
The U-shaped entropy curve in Figure 1 (a) reveals three distinct regions, where each region might benefit from different computational strategies:
  • Overthinking Region (Easy Problems): High correctness rates coupled with high entropy, indicating the model is uncertain about problems it can solve well. This counterintuitive pattern suggests unnecessary computational overhead and motivates a Easy reasoning strategy, which allocates minimal computational resources.
  • Certainty Region (Normal Problems): Low entropy with optimal performance, representing the sweet spot where the model’s uncertainty aligns with task complexity. These problems benefit from Normal reasoning strategy, which allocates standard computational resources.
  • Capability Limit Region (Hard Problems): High entropy with low accuracy, indicating genuine difficulty where the model struggles. These problems may benefit from Hard reasoning strategy, which allocates less computational resources and prevents getting stuck in reasoning.
This analysis naturally leads to a three-tier adaptive strategy: Easy, Normal, and Hard reasoning modes, each tailored to the computational needs revealed by the overthinking phenomenon. We develop such three inference strategies and present oracle experiments with them in the next section.
4 Oracle Experiments with Three Inference Strategies
Figure 2: Our three inference strategy configurations. ||Max|| refers to the predefined maximum token budget for generation. Full prompts and detailed configurations can be found in the Appendix B.
Strategy Temperature Max Tokens Simplified Prompt Template
Easy 0.5 0.4 $\times$ ||Max|| Direct solving with verification
Normal 0.8 1.0 $\times$ ||Max|| Step-by-step methodical approach
Hard 0.4 0.5 $\times$ ||Max|| Resource-aware strategic reasoning
Figure 3: Plots summarizing task performance (y-axis) and efficiency (x-axis) on Qwen3-4B model with various inference strategies. In all datasets, we find oracle strategy (i.e., the strategy that achieves the correct answer while incurring minimal tokens or the strategy incurring minimal tokens if no strategy leads to correct answer) significantly outperforms any uniform setting.
Inference Strategy Set
We design the inference strategies to address the challenges revealed by our overthinking analysis. For Easy problems, we use lower temperature and reduced tokens to prevent the model from exploring unnecessary solution paths. Normal problems receive standard temperature with full token budget to enable comprehensive reasoning in the optimal region where the model’s uncertainty appropriately matches task complexity. For Hard problems, we apply the lowest temperature with moderate token limit to encourage focused, deterministic reasoning while preventing resource waste on problems at the model’s capability limit. Table 3 summarizes the specific configurations for each reasoning strategy, more details can be found in the Appendix B.
Experimental Setting
To establish the theoretical upper bound of our approach, we conduct an oracle experiment using Qwen3-4B across eight reasoning benchmarks. We evaluate on five mathematical reasoning tasks: GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023), AIME 2024 & 2025, and OlympiadBench (He et al., 2024), along with three out-of-domain benchmarks: Minerva (Lewkowycz et al., 2022b), GPQA (Rein et al., 2024), and MMLU-Pro (Wang et al., 2024). For each input, we generates outputs from each of the three inference strategies, Easy, Normal, and Hard. The Oracle strategy is determined through a two-step process: first, we identify all strategies that yield a correct answer, and from this set, we then select the one with the minimal token consumption. The parameter configurations are set as the same as the fixed strategies in Table 3, and the max token limit is 32K. More visualization details can be found in Appendix E.
Results Figure 3 shows that the oracle consistently dominates fixed strategies across all benchmarks, yielding a 7.2% average accuracy gain over the best fixed baseline. Token allocation adapts with problem difficulty, ranging from 198 tokens on GSM8K to 4,675 on AIME25, highlighting the need for adaptive compute budgeting. These results substantiate the overthinking hypothesis and quantify the benefits of difficulty-aware reasoning in both accuracy and efficiency. The oracle’s Pareto-optimal frontier across benchmarks sets actionable performance targets for DiffAdapt.
Figure 4: Overview of the DiffAdapt framework. Top: (i) Data Generation: We sample multiple responses from the proxy model, and compute statistics to heurstically assign inference strategy. This process yields a training dataset for the probe; (ii) Probe Training: We train lightweight probe which takes model’s hidden states after processing the query and predict its inference strategy label; and (iii) Inference with Probe: where the trained probe dynamically selects appropriate inference strategies (Easy/Normal/Hard).
5 The DiffAdapt Framework
To enable LLMs to adapt their inference process based on question difficulty, we introduce a three-stage framework that leverages the model’s internal representations to adapt inference to question difficulty. Figure 4 illustrates our overall approach, consisting of three sequential stages from data preparation to deployment.
5.1 Stage 1: Data Generation with Proxy Model Sampling
The first stage generates training data for the difficulty classifier through a self-supervised approach. Starting from an unlabeled dataset, we use a proxy model—typically the same LLM—to sample responses and compute its entropy and correctness.
For each problem $x$, we prompt the model to generate complete reasoning steps and final results with a maximum length of 32K tokens. We then compute the model’s uncertainty using the same generation entropy calculation described in Section 3. We perform this process with 10 sampling iterations per problem to ensure robust entropy estimates. This generation entropy serves as a proxy for task complexity. Grounded in the U-shaped entropy pattern in Section 3 and the Pareto-optimal Oracle analysis (Appendix C.1), we adopt the following heuristic labeling rule: Normal if correctness $\geq\,\alpha$ and entropy $\leq\,\beta$; Hard if correctness $\gamma$; Easy otherwise (typically the low-difficulty anomaly with moderate correctness but unexpectedly high uncertainty). The thresholds $\alpha,\beta,\gamma$ are set per model using a simple heuristic informed by the observed entropy–correctness distributions; we optionally perform a light sanity check on a small validation split to ensure label stability. Per-model values are reported in Appendix C.2.
5.2 Stage 2: Probe Training on Hidden States
The second stage trains a lightweight probe using the generated labeled data. As shown in the bottom-left of Figure 4, we extract hidden state representations $h_{L}$ from the last layer after prefilling the question.
A small probe $C$ parameterized by $\bm{\theta}$} learns to predict difficulty levels from these hidden states. The probe is implemented as a simple multi-layer perceptron (MLP):
$d=\text{softmax}(W_{2}\cdot\text{ReLU}(W_{1}h_{L}+b_{1})+b_{2})$
where ${\bm{\theta}}=\{W_{1},W_{2},b_{1},b_{2}\}$ are the learnable parameters, and $d$ represents the predicted difficulty distribution over the three classes (Easy/Normal/Hard).
The classifier parameters are optimized by minimizing cross-entropy loss:
$\mathcal{L}({\bm{\theta}})=-\frac{1}{N}\sum_{i=1}^{N}\log P(d_{i}=y_{i}|h_{L}^{(i)};{\bm{\theta}})$
We keep the base LLM weights frozen, requiring only the training of a small probe network.
5.3 Stage 3: Inference with Adaptive Strategy Execution
During the inference, the trained probe dynamically selects a reasoning strategy based on predicted difficulty, as shown in the bottom-right of Figure 4. The process consists of: Difficulty Prediction: Extract hidden states after prefilling the question and predict difficulty using the trained probe. Strategy Selection: Map predicted difficulty to corresponding reasoning strategy (Easy/Normal/Hard). Adaptive Execution: Apply the selected strategy with appropriate computational budget allocation.
This difficulty prediction step does not interfere with the model’s prefilling and decoding processes, making it compatible with most inference optimization techniques including batching, KV cache, prefix cache, and others (Kwon et al., 2023; Zheng et al., 2024).
(a) In-domain performance across reasoning LLMs model architectures. DiffAdapt consistently outperforms fixed strategies across computational budgets.
(b) Out-of-domain performance on Minerva, MMLU-Pro, and GPQA. DiffAdapt maintains effectiveness under domain shift for reasoning LLMs.
Figure 5: Performance across reasoning LLMs model architectures and domains. The x-axis represents different maximum token limit constraints as a percentage of the full token budget, demonstrating how different strategies perform under varying computational budgets. (a) In-domain: DiffAdapt consistently outperforms fixed strategies. (b) Out-of-domain: Effectiveness maintained under domain shift.
6 Experimental Results
6.1 Experimental Setup
Models. We evaluate our framework on five models: three reasoning LLMs (Qwen3-4B (Yang et al., 2025a), DeepSeek-R1-Qwen-7B (DeepSeek-AI et al., 2025), DeepSeek-R1-Llama-8B (DeepSeek-AI et al., 2025)) and three models trained with length control RL training (Nemotron-1.5B (Liu et al., 2025a), [^1] https://huggingface.co/nvidia/Nemotron-Research-Reasoning-Qwen-1.5B ThinkPrune-7B (Hou et al., 2025)). The probe uses a simple MLP structure, trained for 100 epochs using the AdamW optimizer with learning rate 1e-3.
Training Datasets. The probe training dataset consists of a subset of the DeepMath-103K dataset, with 300 problems sampled per difficulty level. Data labeling is performed through Stage I of our DiffAdapt framework 5 to generate difficulty-based strategy assignment.
Baselines. We compare against fixed-strategy baselines that apply exclusively Easy, Normal, or Hard to all problems, as well as the dynamic early-exit method DEER (Yang et al., 2025b). For DEER, we align its maximum generation length with other methods and use the default think threshold of 0.9. This setup highlights the benefit of the proposed DiffAdapt framework over both static and training-free dynamic approaches.
Evaluation Benchmarks. We evaluate on five mathematical reasoning tasks: GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023), AIME 2024 & 2025, and OlympiadBench (He et al., 2024), along with three out-of-domain benchmarks: Minerva (Lewkowycz et al., 2022b), GPQA (Rein et al., 2024), and MMLU-Pro (Wang et al., 2024).
Experimental Design. To demonstrate the probe’s ability to perform difficulty-adaptive reasoning, we mix benchmarks of different difficulty levels. We present averaged results across benchmarks with varying domains, in-domain (GSM8K, MATH500, AIME24&25, OlympiadBench) and out-of-domain (Minerva, GPQA, MMLU-Pro) to show adaptive performance across the difficulty spectrum. Each experiment is run three times; we report the mean across runs. For evaluation, we follow reliability protocols from Ye et al. (2025) [^2] https://github.com/GAIR-NLP/LIMO,Chen et al. (2025) [^3] https://github.com/IAAR-Shanghai/xVerify. Max token limits: for each model–benchmark pair, we first allow a generous cap (e.g., 32K) to observe the model’s longest output and then set a nearby rounded value as the per-benchmark max_tokens used in plots; see Appendix Table 5 for the finalized values and Appendix C.4 for details.
6.2 Reasoning LLMs Results
To validate the generalizability of our DiffAdapt framework, we conducted comprehensive experiments across different reasoning LLMs model architectures and scales. Figure 5 presents the performance-efficiency trade-offs for three representative models on both in-domain (ID) and out-of-domain (OOD) evaluation datasets.
Performance Analysis
DiffAdapt consistently outperforms fixed strategies across all model architectures and domains. On in-domain tasks (Figure 5(a)), Qwen3-4B shows the largest improvements, particularly at higher token budgets. DeepSeek-R1-Qwen-7B demonstrates stable gains, while DeepSeek-R1-Llama-8B exhibits significant benefits across model families. On out-of-domain evaluation (Figure 5(b)), the framework remains effective: DiffAdapt delivers consistent improvements across models and domains, with gains becoming more pronounced as the maximum token budget increases. By contrast, the training-free dynamic baseline DEER performs comparably to the strongest fixed strategy on in-domain datasets but still lags behind DiffAdapt; under distribution shift (Minerva, MMLU-Pro, GPQA), DEER exhibits limited generalization and larger performance degradation. These results confirm DiffAdapt’s robustness across different architectures, scales, and domains.
Key Findings
The results reveal distinct performance patterns for different strategies. Easy strategies achieve high performance with minimal tokens but plateau quickly. Normal and Hard strategies improve continuously with increased computational budgets. This validates averaging across benchmarks of varying difficulty, as it captures the distinct computational requirements of different problem complexities. DiffAdapt exploits these patterns through adaptive strategy selection, consistently outperforming fixed approaches across all token limits. Compared with DEER, DiffAdapt delivers larger and more stable gains across token budgets and domains, indicating that difficulty-aware strategy selection generalizes better than confidence-based early exit. Models with larger inter-strategy performance differences show greater DiffAdapt improvements, confirming that adaptive selection benefits from strategy diversity.
Figure 6: DiffAdapt orthogonality with Length Control RL methods. Performance analysis across three LC-RL trained models on both ID and OOD datasets.
6.3 Orthogonality with Length Control RL Methods
We evaluate DiffAdapt on models trained with Length Control RL (LC-RL) to demonstrate orthogonality with existing training-based approaches. Figure 6 shows results across two LC-RL models: Nemotron-1.5B, ThinkPrune-7B on both in-domain and out-of-domain datasets.
Key Findings. Observing the Easy, Normal, and Hard strategy curves reveals that Easy strategies achieve superior performance in most settings, as LC-RL training adapts these models to solve problems efficiently with low computational cost. DiffAdapt shows slightly lower performance than Easy strategies under low token limits but achieves state-of-the-art results under high token budgets. This aligns with our design philosophy, enabling LLMs to maintain high performance and efficiency across problems of varying difficulty and different token budgets.
This analysis establishes that DiffAdapt can be effectively combined with existing training-based optimization methods, offering a readily integrable solution that enhances reasoning efficiency without requiring modifications to the underlying training paradigm.
6.4 Computational Efficiency Analysis
We quantify computational efficiency in terms of tokens consumed and latency. For token consumption, we measure the average relative token reduction of DiffAdapt compared to a fixed Normal strategy across eight benchmarks $B$:
$\text{Token Savings}=\frac{1}{|B|}\sum_{b\in B}\frac{\text{Tokens}_{b,\text{Normal}}-\text{Tokens}_{b,\text{Method}}}{\text{Tokens}_{b,\text{Normal}}}\times 100\%.$ (1)
Table 1: Comparing token savings (%) of our method and baseline (DEER). Negative means more tokens than the baseline.
Model DiffAdapt DEER
DS-R1-Qwen-7B 9.7% 53.3-53.3%
Qwen3-4B 22.4% 27.5-27.5%
ThinkPrune-7B 10.1%
Table 2: End-to-end inference time comparison. All methods use vLLM backend.
Method Time (minutes) $\downarrow$
vLLM Baseline 64
   + DEER 57
   + DiffAdapt 10
A higher value for Token Savings indicates greater efficiency, meaning more tokens are saved compared to the Normal strategy. Conversely, a negative value signifies that the method consumed more tokens than the baseline.
Token cost
From Table 1, we observe substantial efficiency gains from DiffAdapt across models: Qwen3-4B achieves 22.4% token savings, while DS-R1-Qwen-7B and ThinkPrune-7B reduce usage by around 10%. In contrast, DEER increases token usage relative to Medium (e.g., 27.5-27.5% on Qwen3-4B and 53.3-53.3% on DS-R1-Qwen-7B). The reason is mechanistic: DEER operates under a fixed token budget and chooses continuation length primarily by confidence/probability, which frequently drives generations to the maximum token cap rather than allocating tokens adaptively by problem difficulty.
Latency
[…]
These efficiency gains translate directly into lower inference cost at comparable accuracy levels.
7 Conclusion
We characterize a overthinking phenomenon in LLMs, U-shaped entropy patterns across multiple architectures. This counterintuitive finding challenges the assumption that more computation always improves reasoning.
Our study offers three takeaways: (1) an empirical characterization of overthinking, with a consistent 22–25% entropy reduction from simple to optimal regions that reveals systematic inefficiency; (2) Oracle analysis suggesting a large potential for difficulty-aware inference strategy selection; and (3) DiffAdapt, a lightweight framework that predicts difficulty from hidden states and selects Easy/Normal/Hard strategies, matching or improving accuracy while reducing tokens by up to 22.4% across five models and eight benchmarks.
DiffAdapt requires no LLM retraining, is compatible with common inference optimizations (e.g., batching, KV/prefix caching), and is orthogonal to length-control RL methods. We presents a simple, lightweight solution to allow adaptive computation allocation for LLM reasoning.
Acknowledgements
We thank the members of NYU ML2 for helpful feedback. This work was supported in part by NYU HPC resources, services, and staff expertise.
Appendix
Appendix A Limitations and Future Work
Labeling via proxy models. Our difficulty labels rely on a proxy model and sampling protocol (10 samples at temperature 0.6). Thresholds ($\alpha,\beta,\gamma$) are tuned per model and may require re-calibration when transferring across domains or changing the sampling configuration. In deployment, we recommend a light validation phase to re-establish thresholds.
Failure modes under tight budgets. For particularly hard or error-prone cases, aggressive budget reduction can harm accuracy. A practical fail-safe is to fall back to the Normal strategy when the probe confidence is low, the prefill signal is out-of-distribution, or the selected strategy underperforms recent history.
Probe input sufficiency. Our probe uses hidden states from the question prefill only, which may not fully capture difficulty signals. A promising direction is to incorporate generation-time latent information (e.g., step-wise entropy/logits, attention/KV statistics, or early decoding features) as additional inputs. Such intrusive signals, however, may interrupt standard generation pipelines and introduce extra latency; designing low-overhead hooks is left for future work.
Appendix B Complete Reasoning Strategy
This section provides the complete reasoning strategy configurations used in our three-tier adaptive reasoning framework. Each strategy employs specific prompts designed to guide the model’s reasoning behavior according to the computational requirements identified through our overthinking analysis.
B.1 Easy Strategy Prompt
For problems identified as easy (overthinking region), we use a direct approach to minimize unnecessary computational overhead:
<think>
This looks straightforward. Let me solve it directly while double-checking my approach.
</think>
Configuration:
  • Temperature: 0.5 (lower randomness for direct solving)
  • Max Tokens: 0.4 $\times$ ||Max|| (reduced computational budget)
  • Approach: Direct problem-solving with minimal intermediate steps
B.2 Normal Strategy Prompt
For problems in the optimal region, we employ standard methodical reasoning:
<think>
I’ll break this down into clear, logical steps and solve methodically.
Configuration:
  • Temperature: 0.8 (standard exploration level)
  • Max Tokens: 1.0 $\times$ ||Max|| (full computational budget)
  • Top-p: 0.95 (diverse sampling for comprehensive reasoning)
  • Approach: Step-by-step logical decomposition
B.3 Hard Strategy Prompt
For problems at the capability limit, we focus on efficient resource utilization:
<think>
This appears intricate. I’ll outline the main method while being mindful of computational resources.
Configuration:
  • Temperature: 0.4 (lowest randomness for focused reasoning)
  • Max Tokens: 0.5 $\times$ ||Max||
  • Approach: Strategic method outline to prevent unproductive reasoning loops
B.4 Design Rationale
The prompt design reflects our empirical findings from the overthinking analysis:
  • Easy strategy discourages overthinking by emphasizing directness and verification rather than extensive exploration
  • Normal strategy encourages systematic reasoning with full computational resources for optimal performance
  • Hard strategy promotes efficient strategy selection to avoid getting stuck in computational dead ends
For the hyperparameter choices like temperature and max tokens, we run the grid search to find the best configuration for each strategy.
These strategies, combined with the corresponding sampling parameters, implement the adaptive computational allocation strategy motivated by our U-shaped entropy curve analysis.
Appendix C Additional Experimental Details
C.1 Oracle Experiment Detailed Results
This subsection provides the complete numerical results from our Oracle experiment across eight reasoning benchmarks. Table 3 shows the accuracy and average token consumption for each strategy on every benchmark.
Table 3: Detailed Oracle Experiment Results Across Reasoning Benchmarks With Qwen3-4B
Benchmark Easy Strategy Normal Strategy Hard Strategy Oracle Selection
Acc (%) Tokens Acc (%) Tokens Acc (%) Tokens Acc (%) Tokens
GSM8K 89.90 169.93 93.20 561.80 93.50 511.55 96.20 197.96
MATH 82.80 718.33 96.20 2662.06 93.40 2424.69 98.00 1278.90
GPQA 46.46 812.55 50.51 3022.33 47.98 2952.44 70.20 1001.14
MMLU-Pro 60.36 526.14 65.36 2076.16 62.50 1873.14 74.64 690.04
Minerva 46.69 468.74 55.15 2795.14 53.31 2248.22 65.81 866.31
OlympiadBench 50.96 1628.91 73.63 6722.40 68.59 5895.97 76.89 2756.25
AIME 2024 16.67 3504.90 60.00 10672.23 46.67 10166.87 66.67 4429.37
AIME 2025 16.67 2442.83 53.33 13733.47 40.00 10634.77 56.67 4675.00
[…]
C.2 Model-Specific Threshold Values
This subsection provides the specific threshold values used for difficulty classification across different models in our framework. The thresholds $\alpha$ (correctness), $\beta$ (entropy), and $\gamma$ (correctness) are determined empirically for each model to optimize the strategy assignment performance.
Table 4: Model-Specific Threshold Values for Difficulty Classification
Model $\alpha$ (Normal) $\beta$ (Entropy) $\gamma$ (Hard)
DeepSeek-R1-Qwen-7B 0.85 0.35 0.60
DeepSeek-R1-Llama-8B 0.85 0.35 0.60
Qwen3-4B 0.88 0.32 0.65
Threshold Selection. We use a heuristic procedure guided by the entropy–correctness scatter of each model (no exhaustive search). Concretely, we pick $\alpha$ near the knee where high-correctness, low-entropy points concentrate; choose $\beta$ at the elbow separating low vs. high-uncertainty regimes; and set $\gamma$ to capture the reliability drop-off region in correctness. We optionally verify stability with a small validation split. This selection ensures that:
  • Normal threshold ($\alpha$): Captures problems where the model performs consistently well with low uncertainty
  • Entropy threshold ($\beta$): Distinguishes between confident and uncertain predictions
  • Hard threshold ($\gamma$): Identifies problems beyond the model’s reliable capability range
These model-specific thresholds reflect the inherent differences in reasoning capabilities and uncertainty patterns across different architectures and scales.
C.3 Detailed Algorithmic Procedures
This subsection provides the complete algorithmic descriptions for the three main stages of our DiffAdapt framework. These algorithms detail the implementation procedures that correspond to the conceptual framework presented in Section 5.
Algorithm 1 Data Generation with Proxy Model Sampling
1:Input: Set of problems $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, LLM, thresholds $\alpha$, $\beta$, $\gamma$
2:Initialize labeled dataset $\mathcal{D}\leftarrow\emptyset$
3:for all problem $x$ in $\mathcal{X}$ do
4:  Generate $n=10$ complete reasoning sequences with max length 32K
5:  Compute correctness rate $\mathcal{C}(x)$ and average entropy $\bar{H}(x)$
6:  if $\mathcal{C}(x)>\alpha$ and $\bar{H}(x)<\beta$ then
7:   $y_{\text{label}}\leftarrow\text{`Normal'}$
8:  else if $\mathcal{C}(x)<\gamma$ then
9:   $y_{\text{label}}\leftarrow\text{`Hard'}$
10:  else
11:   $y_{\text{label}}\leftarrow\text{`Easy'}$ $\triangleright$ Overthinking cases
12:  end if
13:  Add $(x,y_{\text{label}})$ to $\mathcal{D}$
14:end for
15:Output: Labeled dataset $\mathcal{D}$
C.4 Maximum Token Limits per Model and Benchmark
This subsection reports the maximum token limits used for each model on each benchmark and describes how they were determined. For each model–benchmark pair, we first ran the model under a generous cap (e.g., 32K tokens) to observe its longest response length in a less constrained setting. We then selected a nearby rounded integer as the per-benchmark max_tokens used in our analyses. This procedure standardizes evaluation across tasks and enables percentage-based truncation in Figures 5 and 6.
Procedure example. With a 32K cap, Qwen3-4B produced longest responses of approximately 1,500 tokens on GSM8K and 18,000 tokens on AIME24; we therefore set max_tokens to 1,500 and 18,000 for those benchmarks, respectively. Analogous rounding was applied to all model–benchmark pairs (see Table 5).
Table 5: Maximum token limits (in tokens) per model and benchmark (ID and OOD). Values are rounded from observed maxima under a large cap (e.g., 32K).
Model GSM8K MATH AIME 2024 AIME 2025 OlympiadBench Minerva MMLU-Pro GPQA
Qwen3-4B 1500 12000 18000 18000 15000 3500 3000 4000
DeepSeek-R1-Qwen-7B 500 3000 15000 16000 5500 1750 3000 5500
DeepSeek-R1-Llama-8B 700 3000 14000 14000 5500 1750 1750 3000
Nemotron-1.5B 3500 4000 7000 6000 5500 5000 3500 5000
ThinkPrune-7B 500 3000 15000 14000 5500 1750 2500 4500
Appendix D Additional Overthinking Analysis Across Model Architectures
To demonstrate the universality of the overthinking phenomenon, we present additional overthinking analysis results for two more model architectures: DeepSeek-R1-Distill-Qwen-1.5B and Nemotron-Research-Reasoning-Qwen-1.5B. These results complement the main analysis presented in Section 3 and provide further evidence that the U-shaped entropy pattern is consistent across different model sizes and architectures.
D.1 DeepSeek-R1-Distill-Qwen-1.5B Overthinking Analysis
Figure 7 shows the overthinking analysis for the DeepSeek-R1-Distill-Qwen-1.5B model. Despite being a smaller 1.5B parameter model, it exhibits the same characteristic U-shaped entropy curve:
  • Simple Problems (Difficulty 1-2): High entropy with good correctness, indicating overthinking behavior
  • Certainty Region (Difficulty 3-6): Reduced entropy with maintained performance
  • Difficult Problems (Difficulty 8+): Increased entropy with declining performance
The entropy reduction of 23.3% from simple to optimal regions demonstrates strong overthinking evidence, consistent with our findings across model architectures.
Figure 7: Overthinking phenomenon in DeepSeek-R1-Distill-Qwen-1.5B model showing the characteristic U-shaped entropy pattern across difficulty levels.
D.2 Nemotron-Research-Reasoning-Qwen-1.5B Overthinking Analysis
Figure 8 presents the analysis for Nemotron-Research-Reasoning-Qwen-1.5B, another 1.5B parameter reasoning model. This model shows the most pronounced U-shaped pattern:
  • Simple Problems (Difficulty 1-2): High entropy with strong correctness, showing clear overthinking
  • Certainty Region (Difficulty 3-6): Significantly reduced entropy with peak performance
  • Difficult Problems (Difficulty 8+): Highest entropy with declining accuracy
This model demonstrates a 21.3% entropy reduction from simple to optimal regions, providing additional validation of the overthinking phenomenon across different reasoning architectures.
Figure 8: Overthinking phenomenon in Nemotron-Research-Reasoning-Qwen-1.5B model demonstrating the universal U-shaped entropy pattern.
Appendix E Additional Oracle Analysis Results
This section presents comprehensive Oracle analysis results across multiple model architectures to validate the generalizability of our findings. We conduct the same Oracle experiment described in Section 3 on three additional models: DS-Qwen-7B, Nemotron-1.5B, and DeepSeek-R1-Llama-8B. These models represent different scales, architectures, and training methodologies, providing robust evidence for the universal applicability of adaptive reasoning strategies.
E.1 DeepSeek-R1-Qwen-7B Oracle Analysis
Figure 9 shows the performance-token trade-offs for DeepSeek-R1-Qwen-7B across all eight reasoning benchmarks. The results demonstrate consistent Oracle superiority with an average improvement of +12.3% over the best fixed strategy, validating our findings across larger model scales.
Figure 9: DeepSeek-R1-Qwen-7B Oracle Analysis: Performance vs. Token Consumption Trade-offs. The Oracle strategy (gold stars) consistently outperforms all fixed strategies across mathematical reasoning tasks (GSM8K, MATH), competition problems (AIME24/25, OlympiadBench), and out-of-domain benchmarks (GPQA, Minerva, MMLU-Pro), achieving optimal Pareto efficiency with an average +12.3% accuracy improvement.
E.2 Nemotron-1.5B Oracle Analysis
Figure 10 presents the Oracle analysis for Nemotron-1.5B, demonstrating that adaptive strategy selection benefits extend to smaller model scales. Despite the reduced parameter count, the Oracle achieves +7.9% average improvement while maintaining superior token efficiency.
Figure 10: Nemotron-1.5B Oracle Analysis: Performance vs. Token Consumption Trade-offs. Even at smaller scale (1.5B parameters), the Oracle strategy demonstrates consistent advantages across all benchmarks, achieving +7.9% average accuracy improvement with efficient token utilization, confirming the scalability of adaptive reasoning approaches.
E.3 DeepSeek-R1-Llama-8B Oracle Analysis
Figure 11 shows the most compelling results from DeepSeek-R1-Llama-8B, which achieves the highest Oracle benefits with +16.2% average improvement. This model demonstrates exceptional token efficiency, with Oracle strategy consuming significantly fewer tokens while achieving superior performance across all benchmarks.
Figure 11: DeepSeek-R1-Llama-8B Oracle Analysis: Performance vs. Token Consumption Trade-offs. This model shows the strongest Oracle benefits with +16.2% average accuracy improvement and exceptional token efficiency. The Oracle strategy achieves superior performance while consuming 10-35% fewer tokens than fixed strategies, demonstrating optimal resource utilization.
E.4 Cross-Model Oracle Analysis Summary
Table 6 summarizes the Oracle analysis results across all four models, demonstrating the universal effectiveness of adaptive strategy selection.
Table 6: Cross-Model Oracle Analysis Summary
Model Parameters Avg. Accuracy Improvement Dominance Rate Token Efficiency
Qwen3-4B 4B +7.2% 100% Mixed
DS-Qwen-7B 7B +12.3% 100% Moderate
Nemotron-1.5B 1.5B +7.9% 100% High
DeepSeek-R1-Llama-8B 8B +16.2% 100% Excellent
Key Insights:
  • Universal Dominance: Oracle strategy achieves 100% dominance rate across all models and benchmarks
  • Scalable Benefits: Performance improvements scale with model capability, ranging from +7.2% to +16.2%
  • Consistent Token Efficiency: All models show improved resource utilization with adaptive strategy selection
  • Robust Generalization: Benefits span mathematical reasoning, competition problems, and out-of-domain tasks
These comprehensive results provide strong empirical evidence that adaptive reasoning strategies offer universal benefits across diverse model architectures, scales, and problem domains, directly motivating the design and deployment of our DiffAdapt framework.